OpenStudy (anonymous):

True or False: (-5 + n) + (5 + -n) = 0

OpenStudy (anonymous):

yes -5 + 5 = 0 -n + n = 0 this is right I think

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

what about True or False: (1/2)(2n + 2) + -2 = n

OpenStudy (anonymous):

false after you expand you are left with -5n+n+5-n=0 then it becomes -5n+5=0 then take the +5 over you get -5n=-5 divide both sides by -5 n=1

OpenStudy (anonymous):

that was for the first question

OpenStudy (anonymous):

no theres no right side, its just -5n+n+5-n

OpenStudy (anonymous):

its false because if you work it out your n=1 and if you put 1 into that equation you end up with 0

OpenStudy (anonymous):

there has to be a right side because the 1st equation = 0

OpenStudy (anonymous):

I think you're thinking too far ahead into algebra, this problem is simple

OpenStudy (anonymous):

maybe i am but thats how i would have worked it out :)

OpenStudy (anonymous):

I did too at the start:)